# What Do Privacy Advertisements Communicate to Consumers?

### v1

@@ 4.3.1. Privacy Awareness
Of the 30 responses collected, 25 discussed the privacy features or functions that were shown in the advertisement or game. The privacy features shown inspired participants to make specific comments related to privacy, with 14 responses explicitly mentioning the word “privacy” in their replies.
@@
In the Apple ad, there were two features that were displayed: “Ask App Not to Track” and “Protect Mail Activity.” When we asked participants to describe the campaign in their own words, in addition to the general synopsis of the campaign, most Apple participants (A1, A2, A4, A5) included a detailed description of the “Ask App Not to Track” feature. A5 said “the ad is about how … using an Apple device can help protect your own data, because there are lots of other apps that could track and use your data, and through Apple you can say that I don’t want you to track it.” However, there was only one mention of the second feature shown in the ad “Protect Mail Activity.” A1 described the protagonist as “going through and selecting options on her iPhone, asking apps not to track her or protecting her email data.” Participants who did not describe either feature (A3, A6) spoke about the ad in much broader terms, with A3 summarizing the features shown in the ad as “apps or certain settings on the phone to where you keep all your data private and secure.” Most Apple participants (A2, A3, A4, A5, A6) mentioned issues surrounding personal data, data sharing, or selling.
@@
For Samsung, the campaign included five privacy features, being privacy dashboard, location permissions, microphone permissions, camera permissions, and Samsung’s Knox Security. The campaign mostly featured the privacy dashboard and location, microphone, and camera sharing permission controls. All participants in this campaign talked about Samsung’s permission controls in general. The location sharing control feature captured the most attention, with four participants (S2, S3, S5, S6) mentioning it in their response, while the camera sharing control feature was mentioned by two participants (S5, S6), and the microphone sharing control was mentioned by one (S6). S6 summarized the ad as “advertising the privacy options that Samsung gives you. Specifically the option to see which apps are using which features such as camera, microphone, location, and you could fine-grain control how they’re using it.”
@@
Each level of Twitter’s DataDash game ends with a button that directs the user to the privacy setting relevant to that game level. All participants experienced level 1 with some progressing to latter levels. No participant clicked on the “Twitter Setting” button at the end of any level, which would have led them to the corresponding privacy setting for that level. When we asked participants to describe the campaign, we had not yet asked any questions containing the words “privacy” or “security.” Thus, two Twitter participants (T3, T5), had yet to realize the game was related to privacy. They only made the connection between DataDash and Twitter’s privacy practices later in the interview, after the interviewer asked: “Were you aware of Twitter’s “Privacy and Safety” features prior to playing DataDash?” Immediately after the interview was over, T5 said that she did not realize there was a privacy connection until halfway through the interview (with her permission we added this comment to her transcript). For the remaining four participants who noticed DataDash’s connection to privacy, there were only two participants who mentioned privacy features in their description of the game. T4 said she felt the game was trying “to show that there are settings to protect … I guess the purpose of the game was to show that I have more control over my data than I realize.” T6 described the purpose of DataDash as “to learn more about Twitter’s features through like, like a simple game.” Of all groups, Twitter participants mentioned privacy the least, with T1 commenting that “I felt like it was a little bit confusing, and it was hard to understand exactly what the game is representing … It gave me an impression that it can be getting dangerous to be on Twitter with all the different elements of privacy concerns.”
@@
The focal point of the WhatsApp campaign was end-to-end encryption. For example WD3 finished her summary of the ad with this sentence: “in the end we find out that 5.5 billion messages are sent in an unsecure way. But if you use Whatsapp, that will not be the case.” When we asked participants to give a summary of the campaign, encryption as a feature was recognized by nearly all participants in this group. WD6, the only participant who did not discuss encryption in WhatsApp specifically, still mentioned unencrypted messages: “[it] kind of draws that parallel to people sending mail regularly versus just sending unencrypted instant message.” All WhatsApp participants mentioned the security or safety of instant messages when discussing the campaign. For example, W5 mentioned “having messages encrypted versus unencrypted using a delivery pigeon.”
@@
When we asked DuckDuckGo participants to describe the ad, all participants discussed tracking and DuckDuckGo’s privacy-friendly search engine and browser. WD4 explained: “Search engines can see everything you type in, and all the results that yield from them, and DuckDuckGo is a solution for that.” However, only two participants (WD1, WD2) were specific about the no-tracking feature. WD1 explained: “Google keeps track and monitors your search history, and DuckDuckGo doesn’t.” All participants mentioned DuckDuckGo’s privacy protective features generally. WD3 described being able to “perform … similar search queries on [DuckDuckGo] platform, but your privacy won’t be impacted.” WD5 was the only one who realized the ad also mentioned the DuckDuckGo mobile app, saying that “I didn’t know before that there was an app for DuckDuckGo.”
@@
We also asked whether participants were aware of the advertised feature (e.g. “Ask App Not to Track”) or function (e.g. encryption or web tracking) prior to watching or interacting with the campaign, and how they came across it (Q13 & Q13B). 18 out of 30 responses indicated prior awareness, including 7 who were non-users. Participants commonly reported that they became aware of a privacy features’ existence after being alerted by the app upon first use (A1, A3, S1, S3, S6) or later when using an app (A2, S6, T2). S6 mentioned both types of prompts in her answer: “it asks you repeatedly when you install it, and then when you use the app.” Others discovered the features through exploring the settings (A2, S2, S4), while some became aware of the feature or concept as it was prominently displayed in the app or website (T5, WD3 for WhatsApp regarding encryption). Participants also reported learning about new features and concepts as a result of updates or policy changes (A5, S5, T4, WD5 regarding DuckDuckGo web tracking feature). T4 described the types of notifications she received from Twitter: “when they update the terms. It comes up as like a pop up on the app. Or sometimes, if they change the terms fully, they send an email.” Interestingly, participants also reported learning about new features and concepts through their usage of other services and apps. WD5 first becoming aware of encryption through “us[ing] VPN services in the past, and also PGP.” WD6 described learning about what encryption is through seeing “a lot of advertising … on Youtube for VPN like TurboVPN and things along those lines”.
@@ 4.3.2. Participants’ Understanding of Privacy Metaphors
As all of the campaigns use metaphors to explain privacy concepts, we investigated what participants understood about these metaphors and whether they were actually useful in illustrating these concepts. We selected one prominent symbol from each campaign and asked participants to share what they thought it might represent (Q.10). We wanted to see how the selected symbol was perceived by participants for each campaign, and whether participants of the same campaign felt similarly about what the symbol could represent. Overall, most participants’ impressions of these metaphors align with the impressions of the authors of this paper, suggesting that the metaphors are largely effective in communicating to non-experts a similar message as they communicate to experts. However, there were occasions where participants failed to notice the privacy metaphor at all, or had a non-privacy related understanding. Two of the metaphors seemed fairly difficult for our participants to interpret correctly.
When we asked participants in the Apple treatment group about the auctioneer, all of them gave responses related to companies selling personal data. Participants said that the auctioneer either represented data sales (A2, A3, A6) or other parties such as app companies, Internet Service Provider (ISP) or smartphone companies (A1, A4, A5). A1 explained that “what they would are really hoping is that you associate … the auctioneer with other cell phone companies … they’re implying that if you are with another cell phone [company] … that phone company will be able to sell your data.” A5 made similar comments, theorizing that “the phone company here could be Apple, or it could be Samsung.”
@@
We asked participants in the Twitter treatment group about the envelopes containing bones that were collected in level 1 and level 2 of the game and received the most diverse interpretations among the campaigns we studied. Two participants (T1, T5) associated the envelopes with data collection, with T5 saying the envelope represented “pieces of data.” Others felt the envelope could be a tweet (T2), spam (T4), or opening a direct message (T6). T3 had the most unique take, stating that he “kind of thought it was a cat, but maybe there was something deeper with that I don’t know.” (Indeed the co-authors were not in agreement, and one of our co-authors also thought the envelopes looked like cats.) Participants’ understanding about the carrier pigeon in the WhatsApp video were mostly uniform, generally themed around instant messaging or the communication channel used. Some thought the pigeon represented SMS message text (WD1 & WD2) or unsafe messages (WD3, WD5). Others felt the pigeon referred to the communication platform (WD3) or the network (WD4), as well as outdated technology (WD1, WD6).
@@ 4.3.3. Participants’ Perceived Purpose of the Campaign
Using the campaign to advertise privacy services or create a privacy-focused company image was the most common perceived purpose, with 24 responses across all campaigns, including all participants in the Apple, Samsung, and DuckDuckGo groups, citing this as the purpose of the campaigns. A1 described the main purpose of the Apple ad as “to sell iPhones, to present them as the most secure option for protecting your identity and your data from unauthorized use.” Participants also felt companies were using these campaigns to advertise themselves as being privacy focused. S6 explained that “the main purpose of video is to make the statement that Samsung very much cares about your privacy.”
Others also felt that the campaigns were issued by the companies to promote themselves as being better than their competitors. We saw 15 comments regarding this across four campaigns (Apple, Samsung, WhatsApp, and DuckDuckGo). All participants in the WhatsApp group cited this as the reason behind the campaign, with WD3 stating that “I think it was purely to get more users to switch on to WhatsApp, instead of using other probably less safe platforms with instant messaging.” Most participants in this group also viewed the purpose of the DuckDuckGo campaign as differentiating themselves from competition. Five DuckDuckGo participants cited this as the main reason behind the campaign. For example, WD6 said, “The main purpose of the video was to show consumers that there is another option or a search platform if they feel like Google is invading their privacy.”
Participants also picked up on potential education efforts and the informative nature of the campaigns, with 13 responses across all groups commenting on this intent. Four Twitter participants (T1, T2, T4, T6) said that DataDash was trying to inform them of some sort of privacy feature or information about Twitter. T1 said he felt “the main purpose of the game was … intended for education.” T2 echoed T1’s response, but was worried the message might have gotten lost: “I think it was to educate people about how Twitter uses our data, but I’m not quite sure it would get the message across correctly because I think people would be so focused on actually [navigating the game] and getting the objects and scoring the points just to beat it that they might miss the entire point of the game.” Indeed, both T3 and T5 failed to recognize the privacy connection between DataDash and Twitter until the interviewer asked a question regarding Twitter’s privacy and safety features. Instead T3 and T5 thought the purpose of the game was entertainment. T3 described the game as “a diversion … something people would do to pass time” and T5 was “not super sure” about the intent of the game “but it was fun.” The Twitter DataDash website explained that the game was designed “to help you understand our privacy policy a little better…. and learn how to take control of your Twitter experience along the way.”
@@ 4.3.4. Impact of Re-Exposure
Since the campaigns we chose were all released in the year before our study, we also asked whether participants had previously seen the ad or played the game prior to the study (Q.6). Only four participants (A5, WD4 WD5 for WhatsApp, WD2 for DuckDuckGo) had previously seen the campaigns. Of those four, three participants came across the campaign while watching TV or using a streaming service, with one participant (WD2) stating he saw the ad after having “read an article about … the Super Bowl ads.”
Previously exposed participants were able to pinpoint specific details that might be overlooked by others the first time they viewed the campaign during the study. For example, WD4 made a comment about a scene in the WhatsApp campaign, stating that: “it’s really funny how the user that they zoomed in on while texting was actually using iMessage, which to my knowledge, is end-to-end encrypted.” In addition, A5 commented: “it’s also very clever that Apple and its other marketing puts the phone over the person’s face, which is also another way of signaling that Apple is a barrier between what you are doing on the Internet and what people can see about you.” A5, who was recruited via snowball sampling, also proactively reached out after the interview and provided a screenshot of similar Apple billboard marketing she had seen in her neighborhood, showing the phone-over-face imagery. With her permission, we included this in our study data.
For non-previously exposed participants, upon second viewing, they were also able to pick up on more details. After re-watching the DuckDuckGo ad WD6 remarked: “I don’t know if I missed something the first time, but I feel like I saw more of the advertisement than I did the last time.” S5, a Samsung non-user, commented on a new detail she observed when watching the ad again: how Android features seem “more granular in terms of they’re literally giving you a list of things that are tracking you.”
4.4. Actionability
@@
Overall we found that the campaigns were not all that effective at teaching viewers how to use privacy features. When we asked participants to direct the interviewer to the privacy features, they seemed to rely on their prior knowledge of the OS and intuition about how to use the interface rather than information provided in the campaign. Some features, such as Apple’s “Ask App Not to Track,” are presented differently depending on whether the user is opening a newly installed app or searching through settings. Thus, the interface shown in the campaign may not exactly match what a user sees when they try to use the feature. In addition, whether participants could easily find the relevant Twitter privacy settings seemed to have little to do with their experience playing the game.

### v2

@@ 4.3.1. Privacy Awareness
Of the 30 responses collected, 25 discussed at least one of the privacy features or functions that were shown in the advertisement or game. The privacy features shown inspired participants to make specific comments related to privacy, with 14 responses explicitly mentioning the word “privacy” in their replies.
@@
In the Apple ad, two features were displayed: “Ask App Not to Track” and “Protect Mail Activity.” When we asked participants to describe the campaign in their own words, in addition to the general synopsis of the campaign, most Apple participants (A1, A2, A4, A5) included a detailed description of the “Ask App Not to Track” feature. A5 said: “the ad is about how … using an Apple device can help protect your own data, because there are lots of other apps that could track and use your data, and through Apple, you can say that I don’t want you to track it.” However, there was only one mention of the second feature shown in the ad “Protect Mail Activity.” A1 described the protagonist as “going through and selecting options on her iPhone, asking apps not to track her or protecting her email data.” Participants who did not describe either feature (A3, A6) spoke about the ad in much broader terms, with A3 summarizing the features shown in the ad as “apps or certain settings on the phone to where you keep all your data private and secure.” Most Apple participants (A2, A3, A4, A5, A6) mentioned issues surrounding personal data, data sharing, or selling.
@@
For Samsung, the campaign included five privacy features: privacy dashboard, location permissions, microphone permissions, camera permissions, and Samsung’s Knox Security. The campaign mostly featured the privacy dashboard and location, microphone, and camera sharing permission controls. All participants in this campaign talked about Samsung’s permission controls in general. The location sharing control feature captured the most attention, with four participants (S2, S3, S5, S6) mentioning it in their response, while the camera sharing control feature was mentioned by two participants (S5, S6), and the microphone sharing control was mentioned by one (S6). S6 summarized the ad as “advertising the privacy options that Samsung gives you. Specifically, the option to see which apps are using which features such as camera, microphone, location, and you could fine-grain control how they’re using it.”
@@
Each level of Twitter’s DataDash game ends with a button that directs the user to the privacy setting relevant to that game level. All participants experienced level 1 with some progressing to latter levels. No participant clicked on the “Twitter Settings” button at the end of any level, which would have led them to the corresponding privacy setting for that level. When we asked participants to describe the campaign, we had not yet asked any questions containing the words “privacy” or “security.” Thus, two Twitter participants (T3, T5), had yet to realize the game was related to privacy. They only made the connection between DataDash and Twitter’s privacy practices later in the interview, after the interviewer asked: “Were you aware of Twitter’s ‘Privacy and Safety’ features prior to playing DataDash?” Immediately after the interview was over, T5 said that she did not realize there was a privacy connection until halfway through the interview (with her permission we added this comment to her transcript). For the remaining four participants who noticed DataDash’s connection to privacy, there were only two participants who mentioned privacy features in their description of the game. T4 said she felt the game was trying “to show that there are settings to protect … I guess the purpose of the game was to show that I have more control over my data than I realize.” T6 described the purpose of DataDash as “to learn more about Twitter’s features through like, like a simple game.” Of all groups, Twitter participants mentioned privacy the least, with T1 commenting that “I felt like it was a little bit confusing, and it was hard to understand exactly what the game is representing … It gave me an impression that it can be getting dangerous to be on Twitter with all the different elements of privacy concerns.”
@@
The focal point of the WhatsApp campaign was end-to-end encryption. For example WD3 finished her summary of the ad with this sentence: “in the end, we find out that 5.5 billion messages are sent in an unsecure way. But if you use Whatsapp, that will not be the case.” When we asked participants to give a summary of the campaign, encryption as a feature was recognized by nearly all participants in this group. WD6, the only participant who did not discuss encryption in WhatsApp specifically, still mentioned unencrypted messages: “[it] kind of draws that parallel to people sending mail regularly versus just sending unencrypted instant message.” All WhatsApp participants mentioned the security or safety of instant messages when discussing the campaign. For example, W5 mentioned “having messages encrypted versus unencrypted using a delivery pigeon.”
@@
When we asked DuckDuckGo participants to describe the ad, all participants discussed tracking and DuckDuckGo’s privacy-friendly search engine and browser. WD4 explained: “search engines can see everything you type in and all the results that yield from them, and DuckDuckGo is a solution for that.” However, only two participants (WD1, WD2) were specific about the no-tracking feature. WD1 explained: “Google keeps track and monitors your search history, and DuckDuckGo doesn’t.” All participants mentioned DuckDuckGo’s privacy-protective features generally. WD3 described being able to “perform … similar search queries on [DuckDuckGo] platform, but your privacy won’t be impacted.” WD5 was the only one who realized the ad also mentioned the DuckDuckGo mobile app, saying: “I didn’t know before that there was an app for DuckDuckGo.”
@@
We also asked whether participants were aware of the advertised feature (e.g., “Ask App Not to Track”) or function (e.g., encryption or web tracking) prior to watching or interacting with the campaign, and how they came across it (Q.13, Q.13B). 18 out of 30 responses indicated prior awareness, including 7 who were non-users. Participants commonly reported that they became aware of a privacy feature’s existence after being alerted by the app upon first use (A1, A3, S1, S3, S6) or later when using an app (A2, S6, T2). S6 mentioned both types of prompts in her answer: “it asks you repeatedly when you install it, and then when you use the app.” Others discovered the features through exploring the settings (A2, S2, S4), while some became aware of the feature or concept as it was prominently displayed in the app or website (T5, WD3 for WhatsApp regarding encryption). Participants also reported learning about new features and concepts as a result of updates or policy changes (A5, S5, T4, WD5 regarding the DuckDuckGo web tracking feature). T4 described the types of notifications she received from Twitter: “when they update the terms. It comes up as like a pop-up on the app. Or sometimes, if they change the terms fully, they send an email.” Interestingly, participants also reported learning about new features and concepts through their usage of other services and apps. WD5 first became aware of encryption through “us[ing] VPN services in the past, and also PGP.” WD6 described learning about what encryption is through seeing “a lot of advertising … on Youtube for VPN like TurboVPN and things along those lines.”
@@ 4.3.2. Participants’ Understanding of Privacy Metaphors
As all of the campaigns use metaphors to explain privacy concepts, we investigated what participants understood about these metaphors and whether they were useful in illustrating these concepts. We selected one prominent symbol from each campaign and asked participants to share what they thought it might represent (Q.10). We wanted to see how the selected symbol was perceived by participants for each campaign, and whether participants of the same campaign felt similarly about what the symbol could represent. Overall, most participants’ impressions of these metaphors align with the impressions of the authors of this paper, suggesting that the metaphors are largely effective in communicating to non-experts a similar message as they communicate to experts. However, there were occasions where participants failed to notice the privacy metaphor at all or had a non-privacy-related understanding. Two of the metaphors seemed fairly difficult for our participants to interpret correctly.
When we asked participants in the Apple treatment group about the auctioneer, all of them gave responses related to companies selling personal data. Participants said that the auctioneer either represented data sales (A2, A3, A6) or other parties such as Internet Service Providers (ISP), smartphone companies, or app companies (A1, A4, A5). A1 explained that “what they would are really hoping is that you associate … the auctioneer with other cell phone companies … they’re implying that if you are with another cell phone [company] … that phone company will be able to sell your data.” A5 made similar comments, theorizing that “the phone company here could be Apple, or it could be Samsung.”
@@
We asked participants in the Twitter treatment group about the envelopes containing bones that were collected in level 1 and level 2 of the game and received the most diverse interpretations among the campaigns we studied. Two participants (T1, T5) associated the envelopes with data collection, with T5 saying the envelope represented “pieces of data.” Others felt the envelope could be a tweet (T2), spam (T4), or opening a direct message (T6). T3 had the most unique take, stating that he “kind of thought it was a cat, but maybe there was something deeper with that I don’t know.” (Indeed the co-authors were not in agreement, and one of our co-authors also thought the envelopes looked like cats.)
Participants had a mostly uniform understanding of the carrier pigeon in the WhatsApp video, generally themed around instant messaging or the communication channel used. Some thought the pigeon represented SMS message text (WD1, WD2) or unsafe messages (WD3, WD5). Others felt the pigeon referred to the communication platform (WD3) or the network (WD4), as well as outdated technology (WD1, WD6).
@@ 4.3.3. Participants’ Perceived Purpose of the Campaign
Using the campaign to advertise privacy services or create a privacy-focused company image was the most common perceived purpose, with 24 responses across all campaigns, including all participants in the Apple, Samsung, and DuckDuckGo groups, citing this as the purpose of the campaigns. A1 described the main purpose of the Apple ad as “to sell iPhones, to present them as the most secure option for protecting your identity and your data from unauthorized use.” Participants also felt companies were using these campaigns to advertise themselves as being privacy-focused. S6 explained that “the main purpose of video is to make the statement that Samsung very much cares about your privacy.”
Others also felt that the campaigns were issued by the companies to promote themselves as being better than their competitors. We saw 15 comments regarding this across four campaigns (Apple, Samsung, WhatsApp, and DuckDuckGo). All participants in the WhatsApp group cited this as the reason behind the campaign, with WD3 stating: “I think it was purely to get more users to switch on to WhatsApp, instead of using other probably less safe platforms with instant messaging.” Most participants in this group also viewed the purpose of the DuckDuckGo campaign as differentiating themselves from competitors. Five DuckDuckGo participants cited this as the main reason behind the campaign. For example, WD6 said: “the main purpose of the video was to show consumers that there is another option or a search platform if they feel like Google is invading their privacy.”
Participants also picked up on potential education efforts and the informative nature of the campaigns, with 13 responses across all groups commenting on this intent. Four Twitter participants (T1, T2, T4, T6) said that DataDash was trying to inform them of some sort of privacy feature or information about Twitter. T1 said he felt “the main purpose of the game was … intended for education.” T2 echoed T1’s response but was worried the message might have gotten lost: “I think it was to educate people about how Twitter uses our data, but I’m not quite sure it would get the message across correctly because I think people would be so focused on actually [navigating the game] and getting the objects and scoring the points just to beat it that they might miss the entire point of the game.” Indeed, both T3 and T5 failed to recognize the privacy connection between DataDash and Twitter until the interviewer asked a question regarding Twitter’s privacy and safety features. Instead, T3 and T5 thought the purpose of the game was entertainment. T3 described the game as “a diversion … something people would do to pass time” and T5 was “not super sure” about the intent of the game “but it was fun.” The Twitter DataDash website explained that the game was designed “to help you understand our privacy policy a little better … and learn how to take control of your Twitter experience along the way.”
@@ 4.3.4. Impact of Re-Exposure
Since the campaigns we chose were all released in the year before our study, we also asked whether participants had previously seen the ad or played the game before the study (Q.6). Only four participants (A5, WD4 & WD5 for WhatsApp, WD2 for DuckDuckGo) had previously seen the campaigns. Of those four, three participants came across the campaign while watching TV or using a streaming service, with one participant (WD2) stating he saw the ad after having “read an article about … the Super Bowl ads.”
Previously exposed participants were able to pinpoint specific details that might be overlooked by others the first time they viewed the campaign during the study. For example, WD4 commented on a scene in the WhatsApp campaign, stating: “it’s really funny how the user that they zoomed in on while texting was using iMessage, which to my knowledge, is end-to-end encrypted.” In addition, A5 commented: “it’s also very clever that Apple and its other marketing puts the phone over the person’s face, which is also another way of signaling that Apple is a barrier between what you are doing on the Internet and what people can see about you.” A5, who was recruited via snowball sampling, also proactively reached out after the interview and provided a screenshot of similar Apple billboard marketing she had seen in her neighborhood, showing the phone-over-face imagery. With her permission, we included this in our study data.
For non-previously exposed participants, upon second viewing, they were also able to pick up on more details. After re-watching the DuckDuckGo ad, WD6 remarked: “I don’t know if I missed something the first time, but I feel like I saw more of the advertisement than I did the last time.” S5, a Samsung non-user, commented on a new detail she observed when watching the ad again regarding how Android features seem “more granular in terms of they’re literally giving you a list of things that are tracking you.”
$4.$ Actionability
@@
Overall we found that the campaigns were not all that effective at teaching viewers how to use privacy features. When we asked participants to direct the interviewer to the privacy features, they seemed to rely on their prior knowledge of the OS and intuition about how to use the interface rather than the information provided in the campaign. Some features, such as Apple’s “Ask App Not to Track,” are presented differently depending on whether the user is opening a newly installed app or searching through settings. Thus, the interface shown in the campaign may not exactly match what a user sees when they try to use the feature. In addition, whether participants could easily find the relevant Twitter privacy settings seemed to have little to do with their experience playing the game.